### v3

# The Race to Efficiency: A New Perspective on AI Scaling Laws

Abstract
As large-scale AI models expand, training becomes costlier and sustaining progress grows harder. Classical scaling laws (e.g., Kaplan et al. [9], Hoffmann et al.[10]) predict training loss from a static compute budget yet neglect time and efficiency, prompting the question: how can we balance ballooning GPU fleets with rapidly improving hardware and algorithms? We introduce the relative-loss equation, a time- and efficiency-aware framework that extends classical AI scaling laws. Our model shows that, without ongoing efficiency gains, advanced performance could demand millennia of training or unrealistically large GPU fleets. However, near-exponential progress remains achievable if the “efficiency-doubling rate” parallels Moore’s Law. By formalizing this race to efficiency, we offer a quantitative roadmap for balancing front-loaded GPU investments with incremental improvements across the AI stack. Empirical trends suggest that sustained efficiency gains can push AI scaling well into the coming decade, providing a new perspective on the diminishing returns inherent in classical scaling.
[…]
The Race to Efficiency: A New Perspective on AI Scaling Laws
Chien-Ping Lu
cplu@nimbyss.com
1 Introduction
The future trajectory of AI scaling is widely debated: some claim that ever-growing models and datasets are nearing practical and theoretical limits [1, 2, 3], while others maintain that ongoing innovations will continue driving exponential growth [4, 5, 6]. For organizations weighing these divergent views, a central question arises: should they “front-load” GPU capacity—relying on the predictable (yet potentially plateauing) gains promised by static scaling laws—or invest in R&D for (possibly unpredictable and hard-to-measure) efficiency breakthroughs, model innovations, and future hardware enhancements? Ultimately, if diminishing returns do indeed loom, how severe might they be in terms of both time and hardware capacity (see Table 2 for an illustrative range of outcomes)?
To address this conceptual gap, we note that any truly enduring “exponential” trend hinges on improving an efficiency metric that reflects both the outcomes and the costs (time, energy, etc.). Historically, Moore’s Law embodied such progress by showing that transistor count per unit area could approximately double every two years [7], while Dennard Scaling [8] kept power usage in check. Turning to AI, classical scaling laws quantify how training loss predictably decreases with increasing compute, provided balanced model, data, and training configurations—referred to as the compute-optimal condition [9, 10]. However, these laws are inherently static: they do not account for the severity of diminishing returns or specify how quickly efficiency must improve to offset these trends over time.
Key Idea: Making Scaling Time- and Efficiency-Aware.
Classical scaling laws [9, 10] posit that $L_{0}\propto C_{0}^{-\kappa}$ for a given static compute budget $C_{0}$. We extend this snapshot into a time- and efficiency-aware framework. Let $L_{0}$ represent the “baseline” loss associated with an initial compute budget. If $\gamma$ denotes the annual efficiency-doubling rate (in $\text{yr}^{-1}$), reminiscent of the 0.5 times per year doubling of transistor density in Moore’s Law, we derive a relative-loss equation that captures how loss evolves over time:
$L(t)\;=\;L_{0}\,R(t),\quad R(t)\;=\;\left(1+\frac{2^{\gamma t}-1}{\gamma\ln(2)\,\times\,1\,\text{yr}}\right)^{-\kappa}.$ (1.1)
Figure 1: AI Scaling and Moore’s Law with Efficiency-Doubling Rates. This plot compares a hypothetical Moore’s Law curve (dashed) with $\kappa=0.4$ and $\gamma=0.5$, against AI scaling curves (solid) at $\kappa=0.048$ (typical of large language models) for various efficiency-doubling rates $\gamma\in\{0,0.5,1,2,3\}$. The horizontal line $R(t)=0.68$ corresponds to a token-prediction probability of 50%, assuming $L_{0}=1.0$. Increasing $\gamma$ drastically reduces the time to cross this threshold. The x-axis represents Time (years), and the y-axis represents Relative Loss $R(t)$. Distinct colors are used for different $\gamma$ values to highlight the impact of efficiency improvements.
Here, $L(t)$ represents the training loss at time $t$ (in years), $L_{0}$ is the initial loss, $R(t)$ is the relative loss, and $\kappa$ is the unitless scaling exponent. The equation captures how training loss evolves over time as efficiency improves. Even with diminishing returns ($\kappa\ll 1$), rapid efficiency gains ($\gamma>0$) can sustain near-exponential progress in AI scaling.
Figure 1 illustrates the interplay between $\kappa$ and $\gamma$. A small $\kappa=0.048$ (typical of large language models) causes the AI scaling curves to flatten significantly over time. With $\gamma=0.5$ (efficiency doubling every two years), reducing $R(t)$ to 0.68 might require approximately 20 years. Increasing $\gamma$ to 2.0 compresses this timeline to well under a decade, while $\gamma=3$ shrinks it further. This demonstrates how higher efficiency-doubling rates can effectively mitigate the limitations imposed by a small $\kappa$.
By contrast, the $\gamma=0$ (flat) curve underscores the severity of diminishing returns, indicating that reaching $R(t)=0.68$ could demand 3,000$\times$ the current GPU capacity or training time—scenarios far beyond real-world feasibility (see Table 2).
Organization.
Section 2 reviews related work on AI scaling. Section 3 formalizes the time-based extension to scaling laws and derives the relative-loss equation. In Section 4, we examine the scaling behavior predicted by the relative-loss equation. Section 5 presents a case study comparing “front-loading” GPUs with sustained efficiency improvements and discusses broader implications. Finally, Sections 6 and 7 summarize key findings and propose directions for empirical validation and future research.
2 Related Work
The study of AI scaling laws has become a cornerstone in understanding how training loss decreases as compute increases under optimized configurations. Kaplan et al. [9] introduced the concept of compute-optimal scaling, demonstrating predictable relationships among model size, dataset size, and compute. Brown et al. [11] reinforced these findings through the scaling behavior of Large Language Models (LLMs) such as GPT-3. Hoffmann et al. [10] refined the framework in the Chinchilla setting, underscoring the importance of balancing model size and dataset size to achieve compute-optimality. Collectively, these foundational studies provide empirical measurements of scaling exponents and form the basis for much of the work in this domain.
Building on these foundations, recent research has explored additional factors influencing scaling laws. Sardana et al. [12] incorporated inference-time compute costs, proposing methods in which smaller models—trained with much larger (potentially synthetic) datasets—can balance efficiency across both training and deployment phases. Snell et al. [13] investigated strategies for optimizing compute specifically at test time. To address various optimizations, Clark et al. [14] introduced sparsity-aware scaling laws for Mixture-of-Experts (MoE) architectures, formalizing an “effective model size.” Building on that framework, Kumar et al. [15] examined precision-aware scaling, showing how precision influences effective parameter counts in a compute-optimal regime.
Despite these advancements, most studies treat compute as a static input rather than a dynamic, time-evolving resource. This paper addresses that gap by integrating empirically established scaling exponents with the temporal dynamics of efficiency improvements, inspired by Moore’s Law [7] and Dennard Scaling [8]. Our work bridges the gap between classical scaling laws and the real-world constraints of time and efficiency, providing a framework for understanding how diminishing returns can be offset by continuous innovation.
3 Mathematical Foundation
We now formalize how to extend classical, static AI scaling laws into a dynamic, time-dependent framework. In particular, we derive a relative-loss equation that unifies traditional loss–compute relationships with a “Moore’s Law-like” perspective on efficiency gains.
3.1 Key Parameters and Notation
Table 1 summarizes the main parameters and variables. In brief, we measure:
    Logical (Model) FLOPs, defined to remain stable, vendor-agnostic, and consistent with industry standards of comparing ”teraflops” among different precisions such as FP16, BF16, FP8, and FP4;
    Power and time, representing real-world usage averaged over a suitable timescale (mean-field assumption);
    A scaling exponent $\kappa$, which captures how loss decreases with total compute, and an efficiency-doubling rate $\gamma$, quantifying how rapidly “usable compute” can grow per unit time and power.
Symbol Definition Units
$t$ Elapsed time since start of training yr
$E(t)$ Compute efficiency at time $t$ PFLOP/yr/MW
$E_{0}$ Baseline/initial efficiency PFLOP/yr/MW
$\gamma$ Annual efficiency-doubling rate yr-1
$P(t)$ Time-varying power usage MW
$P_{0}$ Mean-field power usage MW
$C(t)$ Cumulative compute up to time $t$ PFLOP
$C_{0}$ Initial cumulative compute (snapshot) PFLOP
$\kappa$ Scaling exponent (unitless)
$L_{0}$ Baseline (initial) training loss nats/token
$L(t)$ Training loss at time $t$ nats/token
$R(t)$ Relative training loss: $L(t)/L_{0}$ (unitless)
[…]
3.2 Continuous Efficiency Gains ($E(t)$)
We model efficiency as a continuously evolving resource, reminiscent of how Moore’s Law once described periodic doubling in transistor density. Concretely, let
[…]
where $\gamma$ denotes the annual rate at which efficiency doubles, and $E_{0}$ is the baseline efficiency at $t=0$. Although real improvements may come in discrete jumps, this continuous approximation is mathematically convenient and mirrors how large-scale phenomena (e.g., population growth) are often modeled as exponentials.
Relation to Power and Hardware.
Efficiency, as used here, is dimensionally Logical $\frac{\text{Logical FLOPs}}{\text{time}\times\text{power}}$. In practice, raising $E(t)$ can come from:
    Better hardware (e.g., next-gen accelerators, lower-precision logic, advanced memory or networking),
    Algorithmic gains (e.g., quantization, expert routing),
    Software optimizations (kernel-level efficiency, distributed training overheads), or
    Any combination of the above.
We simply aggregate all these factors into a single, time-varying $E(t)$.
3.3 Cumulative Compute as an Integral ($C(t)$)
Classical scaling laws treat compute $C$ as a static budget. Here, we let $C(t)$ accumulate over time:
$C(t)\;=\;C_{0}\;+\;\Delta C(t),$
where $C_{0}$ is the initial snapshot of compute (reflecting prior investments), and
[…]
If $P(\tau)$ denotes the power allocated to training, then $E(\tau)\,P(\tau)$ is the instantaneous compute throughput (PFLOP/yr). Integrating from 00 to $t$ yields the total additional compute $\Delta C(t)$ beyond the original $C_{0}$.
Mean-Field Assumption.
Rather than modeling $P(\tau)$ at every instant, we approximate
$P(\tau)\;\approx\;P_{0},$
the average power over one year. This “mean-field” approach is common in physics (e.g., average particle collisions) and engineering (e.g., duty cycles). Importantly, this assumption represents an upper bound on performance, as any deviations—such as fluctuations in power usage or suboptimal resource allocation—will result in slower progress in reducing training loss. This makes the mean-field assumption not only mathematically convenient but also practically significant, as it provides an optimistic baseline for evaluating the impact of efficiency improvements.
[…]
Because classical scaling laws directly link $C_{0}$ to $L_{0}$, we define
$C_{0}\;=\;E_{0}\cdot P_{0}\times 1\,\text{yr},\quad\Longrightarrow\quad\frac{C_{0}}{E_{0}\,P_{0}}\;=\;1\,\text{yr}.$ (3.2)
Here, $C_{0}$ is simply the compute obtained by running efficiency $E_{0}$ at power $P_{0}$ for one year. Changing hardware details (e.g., front-loading more GPUs) merely rescales $\,(C_{0},\,E_{0},\,P_{0})$, shifting the initial loss $L_{0}$ but preserving the relative shape of $L(t)$. Consequently, the time-extended scaling law’s trajectory remains the same, regardless of the precise cluster schedule or deployment plan.
Practical Significance.
The mean-field assumption represents an upper bound on performance, as any deviations—such as fluctuations in power usage or suboptimal resource allocation—will result in slower progress in reducing training loss. This makes the assumption not only mathematically convenient but also practically significant, as it provides an optimistic baseline for evaluating the impact of efficiency improvements.
Additionally, the mean-field assumption consolidates human R&D cycles and training sessions. Even if a specific training run takes only a few months, the development process—including model design, data preparation, and hardware procurement—often spans a year or more. By adopting a one-year baseline, our framework naturally aligns with these real-world timelines, providing a practical and intuitive timescale for planning and evaluation. This consolidation ensures that the relative-loss equation remains relevant across multiple iterations of model development and deployment.
3.4 Deriving the Relative-Loss Equation
In the static regime, scaling laws state that the training loss $L$ decreases as a power-law of compute, $L\propto C^{-\kappa}$. Introducing time into the compute accumulation $C(t)$ transforms this into a time-varying equation:
$L(t)\;=\;L_{0}\,\Bigl{(}1+\tfrac{\Delta C(t)}{C_{0}}\Bigr{)}^{-\kappa}.$
[…]
$\Delta C(t)\;=\;\frac{E_{0}\,P_{0}}{\gamma\,\ln(2)}\,\bigl{(}2^{\gamma t}-1\bigr{)}.$
Hence,
$L(t)$ $=\;L_{0}\Bigl{(}1+\frac{\Delta C(t)}{C_{0}}\Bigr{)}^{-\kappa}\;=\;L_{0}\left(1+\frac{2^{\gamma t}-1}{\gamma\ln(2)\,\times\,1\,\text{yr}}\right)^{-\kappa},$ (3.3)
which can be rewritten in relative-loss form:
$R(t)\;=\;\frac{L(t)}{L_{0}}\;=\;\left(1+\frac{2^{\gamma t}-1}{\gamma\ln(2)\,\times\,1\,\text{yr}}\right)^{-\kappa}.$ (3.4)
The relative-loss equation captures how a baseline loss $L_{0}$ evolves over time, provided efficiency improves at a rate $\gamma$. As $\gamma$ increases, $R(t)$ declines more rapidly.
Interpretation.
    Static vs. Dynamic. The relative-loss equation extends static scaling laws into a time- and efficiency-aware domain. When efficiency does not improve ($\gamma=0$), the system effectively reverts to “static” scaling. One could, in principle, keep training on the same hardware for a very long time, making $\Delta C$ grow linearly with time.
    Moore’s Law-Like Perspective. By letting “efficiency” double over time (instead of having a single snapshot), the analysis aligns with the historical notion of transistor-density doubling. Here, $\gamma$ denotes how quickly one can “refresh” hardware and/or optimize software.
3.5 Timescale and Cross-Project Scope
One-Year Baseline.
Our derivation adopts a one-year baseline (via the mean-field assumption (3.2)), so $\Delta C(t)/C_{0}$ measures how compute accumulates beyond that one-year mark. In principle, any timescale—weeks or months—could be used, yielding the same curve shape; but one year naturally aligns with budgeting cycles and hardware-refresh periods. Thus, statements like “doubling efficiency every six months” or “it takes five years to reduce loss below 0.68” gain clear operational meaning for R&D planning.
Multi-Year, Cross-Project Context.
Although the equations might appear to describe a single, multi-year training run, organizations typically develop AI systems iteratively across multiple releases—upcycling existing models [16, 17], refining data pipelines, and introducing new hardware. Each iteration effectively raises efficiency ($\gamma>0$), while training loss (e.g., cross-entropy) offers a monotonic yardstick: newer models must aim for lower loss to surpass predecessors. In this sense, the relative-loss equation becomes a multi-project roadmap: every new wave of improvements compounds upon earlier ones, rather than relying on a single, continuous training job.
4 Analysis of Scaling Behaviors
Having established a time-based framework for AI scaling, we now examine how its two principal parameters—the scaling exponent $\kappa$ and the annual efficiency-doubling rate $\gamma$—shape long-term performance.
4.1 Reduction to Classical Scaling Laws at $\gamma=0$
Starting from the relative-loss equation:
$R(t)\,=\,\Bigl{(}1\;+\;\frac{2^{\gamma\,t}-1}{\gamma\ln(2)\,\times\,1\,\text{yr}}\Bigr{)}^{-\kappa},\quad\text{where}\quad L(t)\;=\;L_{0}\,R(t),$
we now set $\gamma=0$. To handle the limit  $\,2^{\gamma\,t}-1\;\to\;\gamma\,t\,\ln(2)$ for small $\gamma\,t$, we recall the first-order expansion  $\,2^{x}\approx 1+x\,\ln(2)$ for $x\to 0$. Thus,
$\tfrac{2^{\gamma\,t}-1}{\gamma\ln(2)\,\times\,1\,\text{yr}}\;\xrightarrow{\gamma\to 0}\;\tfrac{\gamma\,t\,\ln(2)}{\gamma\ln(2)\,\times\,1\,\text{yr}}\;=\;t.$
Hence, at $\,\gamma=0,$
$L(t)\,=\;L_{0}\,\Bigl{(}1\;+\;\frac{t}{1\,\text{yr}}\Bigr{)}^{-\kappa}.$
Interpretation.
When $\gamma=0$ (no time-based efficiency improvements), this outcome reduces to the static-scaling form $L\propto C^{-\kappa}$. However, we now see how running the same hardware and software for an additional time $t$ merely accumulates compute in a linear fashion. As the nearly flat $\gamma=0$ curve in Figure 1 shows, one must either (a) train for an exceedingly long duration or (b) invest in a massive up-front cluster at $t=0$ to further reduce loss. Thus, the original diminishing returns $(L\propto C^{-\kappa})$ are now made explicit in both time ($t$) and space ($L_{0}$), underscoring why progress inevitably stalls without ongoing efficiency gains ($\gamma>0$).
4.2 Asymptotic Behaviors
Recall that
[…]
and hence,
$R(t)\;\propto\;2^{-\kappa\,\gamma\,t}\quad\text{for large }t.$
Since $\kappa\gamma>0$, $R(t)$ declines exponentially as $t\to\infty$, mirroring the vanishing returns one encounters when investing ever more compute. This parallels the leveling-off observed in Figure 1.
For added intuition, consider a hypothetical analogy to historical Moore’s Law: if one estimates an effective $\kappa\approx 0.4$, then a doubling rate of $\gamma=0.5$ (doubling roughly every two years) might suffice to maintain improvements for a surprisingly long time.
By contrast, modern AI scaling laws typically have much smaller $\kappa\approx 0.05$. Achieving equally robust gains within a decade may therefore require $\gamma\geq 2$ (efficiency doubling every six months) or faster. As we increase $\gamma$, we effectively prolong what could be termed the “productivity cycle”—the window in which near-exponential improvements remain viable.
Figure 2: Sensitivity to baseline perturbations. The horizontal axis shows $\tau$ in years, with $\tau=-1\,\text{yr}$ representing a scenario where the baseline effectively vanishes. Even under large deviations, higher $\gamma$ values preserve robust predictions for time-to-target.
4.3 Sensitivity Analysis
Our model assumes a constant average power budget $P_{0}$ under the mean-field assumption (Equation (3.2)). In reality, infrastructure, workloads, and hardware may fluctuate, introducing uncertainty. To quantify how such changes affect predictions, we add a perturbation $\tau$ via
$\frac{C_{0}}{E_{0}\cdot P_{0}}\;=\;1+\tau\,\text{yr}.$
If $y$ is a target relative loss (say, $R(t(\tau))=y$), then near $\tau=0$, the time-to-target $t(\tau)$ is approximately $1/(\gamma\,\ln 2)$, regardless of $y$. This implies a consistent first-order sensitivity across different baselines.
For large $\gamma$, small shifts in effective compute have an even smaller impact on time-to-target, as illustrated in Figure 2. At, e.g., $\gamma=2$ (2×2$\times$2 × efficiency every six months), the time to reach a moderate loss threshold changes only slightly when the baseline is perturbed. For example, $\tau=1$ extends the time-to-target from 5.06 to about 5.78 years—a modest increase for a long-term projection.
[…]
4.4 Efficiency Doubling Rates and Time Horizons
[…]
    Historical Moore’s Law ($\gamma=0.5$): Doubling every two years may suffice for modest goals over long timescales, but it can push more stringent targets (e.g., 0.7 or lower) out to 15–20 years— far beyond most industry planning cycles.
    Modern Demands ($\gamma\geq 2$): Efficiency doublings every 6–12 months ($\gamma=2$ or 3) compress the entire schedule to a handful of years, in line with contemporary AI’s rapid iteration.
These analyses collectively highlight how $\gamma$ and $\kappa$ together determine the feasibility of near-exponential progress. Even small differences in $\gamma$ drastically alter the trajectory, underscoring the importance of continuous innovation.
5 Implications and Case Studies
Having introduced a time- and efficiency-aware perspective on AI scaling, we illustrate its consequences through several thought experiments. First, we examine two theoretically static scenarios ($\gamma=0$), where compute is “unfolded” either in space (front-loading all GPUs simultaneously) or in time (running a fixed-size cluster for millennia). Next, we consider scenarios with positive $\gamma$, exploring how a balance between up-front GPU investment and sustained efficiency gains can shape multi-year outcomes. Please see Table 2 for the numerical results of these experiments.
Illustrative Scenario.
[…]
5.1 How Severe Are the Diminishing Returns in AI Scaling?
A simplified numeric example demonstrates how purely static assumptions ($\gamma=0$) can result in extremely large hardware requirements or extended training durations, primarily due to the small scaling exponent $\kappa$ (e.g., $\kappa=0.048$).
From $L_{0}=1.0$ to $L=0.68$.
Under the static law (i.e., the relative-loss equation at $\gamma=0$), we have:
$\Bigl{(}1+\frac{t}{1\,\text{yr}}\Bigr{)}^{-\kappa}\;=\;\Bigl{(}1+\frac{\Delta C}{C_{0}}\Bigr{)}^{-\kappa}\;=\;0.68,$
which implies $t\approx 3000$ and $\Delta C\approx 3000\times C_{0}$.
Interpretation:
  • Longer Training (unfold in time): If $C_{0}$ denotes running a baseline cluster for one year, reducing loss from $L_{0}=1.0$ to $L=0.68$ under a $\gamma=0$ assumption would require an additional 3,000 years of training on the same hardware.
  • Bigger GPU Fleet (unfold in space): If the loss reduction must be achieved within a single year, $\Delta C\approx 3000\times C_{0}$ implies that the hardware must be scaled by a factor of 3,000. This would push GPU requirements into the 300 million range, consuming power comparable to the electricity use of an entire continent.
Implications: $\gamma>0$ as Computational Necessity.
Classical scaling laws, when taken in a purely static sense ($\gamma=0$), imply that driving the loss from $L_{0}=1.0$ to $L=0.68$ might require a 3,000$\times$ increase in compute (or equivalently, 3,000 years of training on the same hardware). However, this enormous “3,000$\times$ factor” does not directly translate to realistic time and resource constraints. By introducing $\gamma>0$ (the efficiency-doubling rate), our framework extends the predictive power of classical scaling laws to more realistic settings, where hardware refreshes, architecture refinements, and data-pipeline optimizations occur continually. Mathematically, one now has
$L(t)\;=\;L_{0}\,\Bigl{(}1+\frac{2^{\gamma t}-1}{\gamma\ln(2)\,\times\,1\,\text{yr}}\Bigr{)}^{-\kappa},\quad\gamma>0,$
so that periodic improvements in “usable compute” offset the high cost of further reducing loss.
In practice, major AI labs do achieve these ongoing gains via hardware refreshes, architecture refinements, and data-pipeline optimizations. Our framework formalizes this necessity, preserving the predictive power of classical scaling laws while integrating time and efficiency for more realistic outcomes.
Scenario Initial GPUs $\gamma$ $L_{0}$ $R(t)$ Time to $L=0.68$
Unfold in Space $\sim\mathbf{3\times 10^{8}}$ 0 0.68 1.00 $\sim 1$ yr
Unfold in Time 100k 0 1.00 0.68 $\sim\mathbf{3{,}000}$ yrs
Baseline 100k 0.5 1.00 0.68 $\sim 20$ yrs
Turtle 10k 3.0 1.12 0.61 $\sim 5$ yrs
Hare 150k 2.0 0.95 0.71 $\sim 5$ yrs
Table 2: Five illustrative scenarios targeting a training loss of $L=0.68$ on a model scaled 10$\times$ beyond the baseline. Unfold in Space and Unfold in Time both set $\gamma=0$ (strictly static). They differ in whether we pack all needed resources into one year ($3\times 10^{8}$ GPUs) vs. stretching the same 100k-GPU baseline across millennia. By contrast, Baseline, Turtle, and Hare assume $\gamma>0$, meaning efficiency improves continuously rather than relying on a single static “snapshot.”
 To reduce loss below 0.68 with $\gamma=0$ requires exponentially more GPUs, due to the small scaling exponent (diminishing returns).
5.2 A Multi-Year Case: Baseline, Turtle, and Hare
We now consider three illustrative scenarios, each aiming to reduce training loss from an initial loss $L_{0}$ to a final target of $L=0.68$. Although they share the same scaling exponent $\kappa$ and ultimate objective, they differ in:
    Starting Training Loss ($L_{0}$), which depends on how many GPUs were initially allocated, and
    Annual Efficiency-Doubling Rate ($\gamma$).
In particular:
    Baseline: Begins with $100{,}000$ GPUs at $\gamma=0.5$, reflecting a historical rate of doubling roughly every two years.
    Turtle: Starts with fewer GPUs ($10{,}000$) but targets a higher $\gamma=3.0$ (tripling annually) to see if an “extreme efficiency push” can catch up.
    Hare: Begins with $150{,}000$ GPUs (about 1.5×1.5$\times$1.5 × the Baseline’s cluster) at $\gamma=2.0$, balancing stronger up-front capacity with a still-robust annual doubling rate.
Common Setup and Baseline Loss $L_{0}$.
We assume each scenario runs for one initial year, after which the training loss is recorded as a new baseline $L_{0}$. Formally, if $L_{\mathrm{init}}$ is the loss before this year of training, and $\Delta C$ is the additional compute gained in that year (relative to a baseline $C_{0}$), then:
$L_{0}\;=\;L_{\mathrm{init}}\,\bigl{(}1+\frac{\Delta C}{C_{0}}\bigr{)}^{-\kappa}.$
Fewer initial GPUs yield a negative $\Delta C$, as the compute achieved falls short of the baseline $C_{0}$, raising $L_{0}$. Conversely, more GPUs yield a positive $\Delta C$, as the compute achieved exceeds the baseline, lowering $L_{0}$. This relationship underscores the trade-off between initial resource allocation and the resulting training loss.
Turtle vs. Hare in Practice.
A scenario inspired by DeepSeek-V3 [18] could exemplify the Turtle approach:
    Smaller GPU Fleet: Starting with fewer GPUs (e.g., 2,000 rather than 10,000) and aiming for a higher efficiency-doubling rate ($\gamma>2$) to offset the lower initial capacity.
    High-Impact Optimizations: Even on mid-range devices, leveraging advanced hardware features (e.g., FP8) or specialized software (e.g., Mixture-of-Experts) can systematically increase $\gamma$.
In practice, DeepSeek-V3 has achieved superior performance across multiple benchmarks, showcasing how a Turtle-style strategy can deliver both efficiency and state-of-the-art results. Over time, however, growing model or dataset requirements might still necessitate some expansion of the GPU cluster.
By contrast, a Llama 3 (405B)-like [19] scenario may follow the Hare approach:
    Front-Loading a Massive Cluster: Deploying a large fleet (e.g., 30,000 GPUs) to rapidly lower training loss in the first year, reducing initial turnaround but at a high capital cost.
    Selective Hardware and Architectures: Potentially a more conservative stance on certain capabilities (e.g., partial adoption of FP8 or limited use of Mixture-of-Experts), with plans to integrate further optimizations in subsequent iterations.
In short, these contrasting strategies highlight how organizations may balance a large initial GPU investment (Hare) with incremental efficiency gains (Turtle) to manage steep diminishing returns. Regardless of fleet size or GPU type, the unifying principle is continuous efficiency improvement across hardware, software, and data pipelines—driving the sustained innovation ($\gamma$) that underpins modern AI scaling.
5.3 Scaling Laws as the Driving Force for Innovations
Sustaining progress in AI scaling demands a deliberate focus on innovation to accelerate efficiency gains and counteract diminishing returns. Interestingly, this driving force stems from the same principles that underpin classical scaling laws, once extended into a time- and efficiency-aware framework.
Logical Compute as Optimization-Agnostic.
Because scaling laws inherently treat compute in a model-agnostic manner, logical compute is defined as though the model were both dense and full-precision. This prevents conflating compute with efficiency, the latter reflecting tangible gains from optimizations like sparsity or reduced-precision formats [15, 14]. Without this separation, an architecture such as DeepSeek-V3 [18]—which achieves a 17×17$\times$17 × higher real-world efficiency than Llama 3 (405B) by leveraging sparsity (Mixture-of-Experts), FP8 arithmetic, and other refinements (see Appendix A)—would appear artificially “smaller.”
By anchoring “compute” to the model’s full architectural capacity and attributing actual speedups to $\mathrm{Time}\times\mathrm{Power}$, the optimization-agnostic nature of scaling laws remains intact. Researchers can innovate freely—balancing accuracy, power, and training cost—without altering the fundamental compute measure itself. Meanwhile, the annual efficiency-doubling rate $\gamma$ quantifies how rapidly these real-world optimizations accumulate, fueling near-exponential progress over multi-year development cycles.
Cumulative Compute as a Compounding Process.
Classical AI scaling laws (e.g., $L_{0}\propto C_{0}^{-\kappa}$) provide a “snapshot” for training loss under a fixed compute budget. Once we generalize to
$L(t)\;=\;L_{0}\,\Bigl{(}1+\frac{\Delta C(t)}{C_{0}}\Bigr{)}^{-\kappa},$
the additional compute $\Delta C(t)$ builds over time, transforming that static snapshot into a compounding process. Every phase of progress leverages all prior compute investments, allowing near-exponential improvements when efficiency ($\gamma>0$) continues rising. This view clarifies why ongoing innovation is indispensable: each incremental gain can magnify the returns of earlier investments, thereby sustaining AI scaling despite inherently small exponents ($\kappa$).
5.4 Outlook
Connection to Industry Trends.
Recent industry data highlights significant advancements in AI energy efficiency, with NVIDIA reporting a 45,000$\times$ improvement in energy efficiency for AI inference over the past eight years—equating to doubling approximately every six months ($\gamma\approx 2$[20]. This progress is exemplified by the latest GB10, which integrates the Grace CPU and Blackwell GPU into a desktop form factor [21]. Concurrently, organizations like OpenAI, Google, and Meta continue to refine hardware, software, data pipelines, and infrastructure to sustain rapid improvements in efficiency. For example, DeepSeek-V3 [18] achieves substantial efficiency gains through innovations such as FP8 arithmetic, sparsity, and Mixture-of-Experts, demonstrating the potential of architectural optimizations. The success of platforms like Grace Blackwell and innovations such as DeepSeek-V3 underscore the critical importance of prioritizing efficiency improvements in AI development, particularly for organizations with constrained compute resources or sustainability targets. This sustained trajectory of efficiency improvement directly supports the exponential performance gains necessary for training modern large-scale models.
Policy Implications.
By explicitly embedding the efficiency-doubling rate $\gamma$ within AI scaling laws, our framework elevates innovation across the AI stack from an implicit assumption to a measurable driver of progress. Rather than a “compute arms race,” AI scaling becomes a “race to efficiency”: leaders and policymakers can set explicit targets (e.g. doubling efficiency every six months) and synchronize development roadmaps to sustain a high $\gamma$. Much as Moore’s Law once provided concrete milestones for transistor scaling, AI practitioners can now anchor multi-year plans on tangible efficiency benchmarks, fueling the compounding gains crucial for sustained, real-world AI advances.
6 Conclusion
This work presented a time- and efficiency-aware foundation for AI scaling, revealing how a “race to efficiency” naturally emerges once classical, static scaling laws account for ongoing efficiency gains. Three key insights form a mutually reinforcing cycle:
    Time-Extended Perspective. While classical laws link loss to compute at a single snapshot, recognizing “efficiency doubling” over months or years shifts that static view into a dynamic, near-exponential trajectory.
    Efficiency-Centric Focus. Because scaling laws reliably map compute to performance gains, the crucial question becomes how efficiently compute accumulates over time. In this light, efficiency doubling emerges as both a computational and practical necessity to mitigate steep diminishing returns.
    Innovation as Core to Scaling. Once efficiency is central, continual optimizations across the AI stack no longer appear as external “fixes” but as integral parts of the scaling process. These incremental improvements compound over multiple training cycles and product generations, reinforcing the time-extended perspective.
Looking ahead, constraints akin to those that once challenged transistor scaling may ultimately call for new paradigms. Yet the tension between diminishing returns and time-extended efficiency gains will likely remain a defining force in AI’s technological evolution.
7 Limitations and Future Work
While the relative-loss equation offers a unified perspective on AI scaling progress, its practical value and generality warrant further exploration. Below, we highlight several avenues for extending and refining this framework.
Empirical Validation and Transparency
The insights in this paper rest on theoretical constructs and empirically observed scaling exponents. Nevertheless, comprehensive validation against real-world data is crucial. Greater transparency in reporting relative training loss, cumulative compute, and efficiency-doubling rates could enable more rigorous cross-study comparisons. Industry-wide data sharing, standardized benchmarks, and consistent evaluation protocols—akin to those once used for guiding semiconductor progress—would help verify the predictive power of the relative-loss equation. Such efforts could also inform resource-allocation decisions, model architectures, and targeted efficiency improvements.
Generalizing to Multi-Phase Growth
Although our framework focuses on AI training, it could be generalized to capture early, sub-exponential “kick-off” phases or logistic transitions in other domains—ranging from technology adoption to broader economic processes. Such a generalization would provide a unified view of how systems evolve from an initial ramp-up to potentially exponential (or S-curve [22]) growth trajectories.
Optimal Usage vs. Raw Compute
The relative-loss equation assumes compute-optimal usage, where model size, dataset size, and training strategy are balanced to fully exploit available resources. Simply increasing GPUs or peak compute does not guarantee improved performance if scaling principles are not followed.
For instance, training a large model without adequate data or failing to tune hyperparameters may not yield the expected loss reductions. Likewise, imbalanced scaling between model and data can prevent the envisioned gains. The equation thus reflects a best-case trajectory, assuming that each increment in compute efficiency translates directly into effective training progress.
Extending the Framework to Inference and Test-Time Scaling
While our current formulation focuses on training dynamics, extending the relative-loss equation or developing analogous constructs for inference-time scaling [13, 23] could yield a more holistic view of AI system performance. In particular, because large models are increasingly used to generate new training data and perform on-the-fly or iterative refinement, higher inference efficiency can accelerate the training pipeline rather than merely reduce deployment costs.
As a result, understanding how inference-time efficiency improvements translate into faster throughput, lower latency, or expanded data pipelines may be crucial, given that deployment considerations (and the downstream feedback loop into training) increasingly shape large-model design choices. A unified framework that addresses both training and inference could thus clarify how hardware roadmaps, data engineering, and algorithmic optimizations interact across the entire lifecycle of AI systems.
Evolving Concepts of Compute-Optimality
Traditionally, compute-optimal scaling assumes a static dataset and a fixed model configuration. In reality, both datasets and models evolve over time. Approaches such as upcycling pretrained models [24, 16, 17], and dynamically adapting model size or precision introduce new optimization strategies. Similarly, dataset generation and curation—leveraging synthetic data [25] or reasoning-based selection [26]—blur the boundary between model development and data sourcing.
As these strategies mature, future frameworks must treat datasets, models, and compute budgets as interconnected and evolving. Such adaptability ensures the relative-loss equation and similar scaling models remain relevant.
Appendix A Invariance of Logical Compute
This appendix explains why we define logical compute as dense and full-precision and how it underpins a fair comparison—illustrated with DeepSeek-V3 and Llama 3 (405B)—even when actual GPU hours differ across hardware or optimization strategies.
Why $\kappa$ Remains Unchanged by Optimizations
Suppose the training loss depends on model size $N$ and dataset size $D$ as
$L(N,D)\;=\;A\,N^{-\alpha}\;+\;B\,D^{-\beta}\;+\;E,$
where $\alpha,\beta>0$. For a fixed compute budget $C$, the compute-optimal pairs $(N^{*},D^{*})$ satisfy
$N^{*}\;\propto\;C^{\tfrac{\beta}{\alpha+\beta}},\quad D^{*}\;\propto\;C^{\tfrac{\alpha}{\alpha+\beta}},$
so substituting back yields
$L(C)\;\propto\;C^{-\tfrac{\alpha}{\alpha+\beta}}\;=\;C^{-\kappa},\quad\kappa\;=\;\frac{\alpha}{\alpha+\beta}.$
Hence, $\kappa$ depends only on $\alpha,\beta,$ not on how we optimize model parameters or numeric formats. In other words, whether one uses low-precision arithmetic, sparsity, or a Mixture-of-Experts design, the fundamental exponent $\kappa$ stays invariant.
Defining Logical Compute
We define logical compute as though the model is both dense and full-precision. Specifically:
[…]
where
  • $N$ = total model parameters,
  • $D$ = total training tokens,
  • The factor 6 accounts for forward/backward passes and parameter updates.
This ensures that “compute” consistently reflects the full architecture, independent of sparsity or precision. Actual speedups (e.g., from FP8 or MoE) appear separately in reduced $\mathrm{Time}\times\mathrm{Power}$, rather than shrinking the fundamental compute measure.
Case Study: DeepSeek-V3 vs. Llama 3 (405B)
To illustrate why logical compute is kept dense, consider two models that achieve broadly similar large-scale outcomes yet differ in real-world GPU usage:
  • DeepSeek-V3 [18]:
    • $N\approx 671\,\mathrm{B}$ (parameters),
    • $D\approx 14.8\,\mathrm{T}$ (training tokens),
    • Real GPU usage: $\sim 2.78\,\mathrm{M}$ GPU-hours on H800, adjusted to $\sim 2.224\,\mathrm{M}$ GPU-hours at H100 equivalence.
  • Llama 3 (405B) [19]:
    • $N\approx 405\,\mathrm{B}$,
    • $D\approx 2.0\,\mathrm{T}$,
    • Real GPU usage: $\sim 30.84\,\mathrm{M}$ GPU-hours on H100.
Summarizing the Computations.
Model Parameters (B) Data Tokens (T) Logical Compute (PFLOPs) GPU Hours (M, H100-Eq.) Relative Efficiency
[…]
Llama 3 (405B) 405 2.0 $4.86\times 10^{15}$ 30.84 1.0 (baseline)
Table 3: DeepSeek-V3 vs. Llama 3 (405B). Both achieve large-scale performance but differ in GPU hours. Logical compute (assuming dense, full-precision) is high in both cases, yet DeepSeek-V3 real-world efficiency is about 17×17$\times$17 × Llama 3’s.
As shown in Table 3, logical compute for DeepSeek-V3 is slightly larger than Llama 3, reflecting its extra parameters and bigger training set. However, real-world GPU hours for DeepSeek-V3 are far lower than Llama 3, thanks to hardware/software optimizations (e.g., sparsity, FP8). Defining compute in a purely “dense” sense prevents conflating those optimizations with the model’s intrinsic size.
Relative Efficiency.
One can define a “relative efficiency” factor:
Relative Efficiency=Logical Compute (FLOPs)GPU-Hours (H100 $\text{Relative Efficiency}\;=\;\frac{\text{Logical Compute (FLOPs)}}{\text{GPU-Hours (H100 Eq.)}},$
then normalize Llama 3 (405B) at 1.0. Under that measure, Table 3 shows DeepSeek-V3 is about 17×17$\times$17 × more efficient. Crucially, this ratio does not shrink its logical compute (it remains at  $\,5.95\times 10^{15}\,\mathrm{PFLOPs}$), but records gains via fewer GPU hours.
Counterarguments and Responses
Should We Adjust Logical Compute for Sparsity or Precision?
Some suggest reducing FLOPs to reflect only the fraction of parameters activated per token (experts-per-token in MoE) or the lower numeric cost (e.g., from FP8). However, that merges two concepts:
  • Full Model Complexity: The entire parameter space at full precision, representing the model’s theoretical capacity.
  • Real-World Efficiency Gains: Achieved by using only a subset of parameters, or fewer bits per operation, thus lowering time and power.
Conflating them penalizes architectures that are inherently more efficient (like Mixture-of-Experts). Instead, we keep the definition of “logical compute” dense and record any real-world speedups in the denominator ($\mathrm{Time}\times\mathrm{Power}$). This way, a model reaps the benefits of advanced routing or quantization ($\gamma>0$) without artificially reducing its fundamental FLOP count.
Is “Effective Model Size” More Accurate?
Although some frameworks define an effective size $N_{\mathrm{eff}}$ for MoE [14] or reduced precision [15], such an approach can hide the full parameter space. Not all parameters are active per token, but they still exist, providing capacity for generalization and future scaling. By keeping “logical compute” dense, we preserve fairness across architectures. If a model invests in sophisticated routing or quantization, the advantage should appear as a lower $\mathrm{Time}\times\mathrm{Power}$, not by discarding parameters from the total.
Summary
Because $\kappa$ depends solely on the power-law slope linking model size, dataset size, and compute, it is invariant to whether a model employs sparsity, lower-precision arithmetic, or mixture-of-experts routing. Defining logical compute as though it were both dense and full-precision provides a consistent baseline for comparing very different architectures. Real-world efficiency gains, meanwhile, show up in $\mathrm{Time}\times\mathrm{Power}$, thereby highlighting the genuine speedups achieved by hardware or software improvements. This framework allows us to preserve the foundational exponent $\kappa$ from classical scaling laws while giving proper credit for engineering advances in accelerators, memory systems, or training algorithms.